### v1

# Epitome: Pioneering an Experimental Platform for AI-Social Science Integration

Abstract
The integration of Large Language Models (LLMs) into social science experiments represents a transformative approach to understanding human-AI interactions and their societal impacts. We introduce Epitome, the world’s first open experimental platform dedicated to the deep integration of artificial intelligence and social science. Rooted in theoretical foundations from management, communication studies, sociology, psychology, and ethics,Epitome focuses on the interactive impacts of AI on individuals, organizations, and society during its real-world deployment. It constructs a theoretical support system through cross-disciplinary experiments. The platform offers a one-stop comprehensive experimental solution spanning "foundation models—complex application development—user feedback" through seven core modules, while embedding the classical "control-comparison-comparative causal logic" of social science experiments into multilevel human-computer interaction environments, including dialogues, group chats, and multi-agent virtual scenarios. With its canvas-style, user-friendly interface, Epitome enables researchers to easily design and run complex experimental scenarios, facilitating systematic investigations into the social impacts of AI and exploration of integrated solutions.To demonstrate its capabilities, we replicated three seminal social science experiments involving LLMs, showcasing Epitome’s potential to streamline complex experimental designs and produce robust results, suitable for publishing in the top selective journals. Our findings highlight the platform’s utility in enhancing the efficiency and quality of human-AI interactions, providing valuable insights into the societal implications of AI technologies. Epitome thus offers a powerful tool for advancing interdisciplinary research at the intersection of AI and social science, with potential applications in policy-making, education, and industry, as well as in customer service, financial advising, consulting, manufacturing design, and more.
Keywords AI  $\cdot$ LLMs  $\cdot$ Social Science Experiment Platform  $\cdot$ Human-AI Interaction
Significance Statement
Epitome,the world’s first open experimental platform dedicated to the deep integration of artificial intelligence and social science. Epitome offers an user-friendly interface enabling researchers to design complex human-AI interaction experiments—such as AI-human cooperation, human evaluation on AI’s sociality, and AI ethical dilemma studies etc, by offering a one-stop comprehensive experimental solution in multilevel human-computer interaction environments. By replicating seminal studies on AI’s productivity and creativity impacts, we demonstrate its capacity to automate data collection, ensure reproducibility, and scale interdisciplinary research. Epitome addresses critical gaps in evaluating AI’s societal effects, offering policymakers, educators, and industries a robust tool to ethically optimize AI deployment while advancing theoretical frameworks for human-AI collaboration. To help users familiarize themselves with the platform, we have provided a detailed introduction to the functionalities of Epitome on GitHub (https://github.com/epitome-AISS/epitome).
Introduction
The convergence of artificial intelligence and social science creates unprecedented opportunities for understanding human behavior and social dynamics in the digital age [1]. This convergence manifests in two distinct research streams: ’AI for Social Science,’ which employs AI tools to enhance social science methodology, and ’Social Science of AI,’ which examines how AI systems influence individuals and societies. These interdisciplinary domains signify a fusion of technological advancements and social investigations, utilizing AI tools to dissect intricate social phenomena [2, 3, 4, 5], while also applying social science methodologies to grasp the societal implications of AI [6, 7, 8]. Despite this promising intersection, researchers face significant methodological barriers when designing experiments that either utilize AI capabilities or investigate human-AI interactions. We introduce Epitome, an experimental platform that bridges this methodological gap through integrated tools fordesigning, implementing and analyzing social science experiments involving large language models.
Experimental studies play a pivotal role in integrating AI and social science by enabling hypothesis testing, variable manipulation, and the observation of causal relationships, offering critical insights into human behavior and AI interactions. In ’AI for Social Science’, experiments have advanced the simulation of subjects [9], creation of social environments [10], automation of processes [11], and prediction of social outcomes [12], deepening our understanding of social dynamics. In "Social Science of AI," experiments have been key to assessing AI’s capabilities in personality and socioeconomic status [13], psychological cognition [14], decision-making [15], and societal impacts such as labor market crises [7], collaboration effectiveness [16, 17], and trust-building [18]. These studies inform policy and enhance public awareness, driving progress in both fields.
These studies inform policy development and public discourse while advancing both fields. However, three obstacles impede further progress. First, technical barriers prevent many social scientists from effectively incorporating AI tools into their research [19]. Second, conventional methodological frameworks inadequately address the complexities of human-AI interaction. Third, experimental designs involving AI often require programming expertise beyond the training of most social scientists. Large language models offer a solution by automating complex technical tasks and enabling more sophisticated experimental designs without requiring specialized coding skills. These models make experimental research more efficient. They let researchers test hypotheses, vary variables, and track causal links [20]. AI can precisely control complex social interactions. This control enables a new paradigm called Integrative Experiments [21]. The paradigm stresses constant testing and refinement of theories. Realizing this paradigm requires a robust experimental platform.
The lack of insights from the social sciences may leave AI researchers struggling to understand the broader social contexts in which their technologies operate, potentially leading to misapplications. This gap can stifle innovation, as social theories often serve as wells of creative problem-solving approaches. Additionally, the ability to assess the impacts of AI technologies may be compromised, resulting in one-off evaluations that fail to capture long-term social consequences. Current safety evaluations are overwhelmingly capability-oriented—approximately 85.6% focus solely on model outputs—leaving human-interaction and systemic-impact assessments largely unexplored [22]. For instance, the overreliance on standardized datasets in current AI evaluations could hinder the development of models that navigate real-world social interactions [23, 24]. While these datasets are crucial training resources, they often lack the dynamic and nuanced qualities inherent in human-to-human interactions [25, 26]. Human feedback, despite its susceptibility to biases and inefficiencies, is vital for capturing the subtleties of social behavior that are difficult to encode in static data [27, 28]. This need becomes particularly evident when considering the iterative, context-dependent nature of human socio-cognitive learning, which frequently draws from real-world experiences [29] and embraces diverse learning modes [30, 31]. Understanding social theories is crucial for addressing ethical issues in AI deployment. Without this understanding, AI researchers may overlook important societal impacts of their work.
Additionally, the inclination of AI researchers to prioritize automation [11] and dehumanization [32] in simulated environments may hinder the development of AI agents capable of engaging in meaningful social interactions. While simulated environments offer controlled settings for experimentation, they often fail to replicate the complexity and richness of real-world social contexts [33]. Such attempts may yield limited results. Numerous agent interaction experiments currently lack guidance from social science theories and remain at the level of descriptive analysis. To truly understand and replicate human social behavior, AI models must be grounded in a deep understanding of human psychology and sociology [34].
As per existing literature, a standard Social Science of AI (SS of AI ) or AI for Social Science(AI for SS) experiment typically comprises 12 tasks, organized across four distinct stages: Experiment Proposal, Experiment Preparation, Experiment Execution, and Data Collection & Analysis . Researchers typically undertake these 12 tasks individually, which include designing the experiment, enlisting engineers for interface development, orchestrating participant recruitment, and overseeing data collection. This process demands a significant investment of time, labor, and resources [19]. Moreover, the complexity of such studies requires the use of advanced tools and workflows, including automated participant management systems and interactive visualization panels, to effectively handle the intricacies of these endeavors(for details, see Supplementary Information SI.1).
To address these challenges, and recognizing the lack of existing platform tools capable of resolving them, we introduce Epitome, an innovative experimental platform designed for the intersection of social science and AI research. Epitome introduces three methodological innovations to social science experimentation. First, it enables dynamic experimental designs where multiple AI agents interact with human participants, allowing researchers to precisely control and manipulate variables in complex social scenarios. Second, it implements standardized measurement of human-AI interactions through automated data collection and analysis protocols that preserve ecological validity. Third, it facilitates experimental reproducibility by packaging experimental designs, stimuli, intervention parameters and analysis workflows into shareable templates. These innovations allow researchers to conduct experiments that were previously impractical or impossible due to resource constraints or technical barriers.
Result
Epitome:Innovative Methodologies for Integrating AI and Social Science
Layout and Key Functions By constructing $55$ layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com [^1] http://www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
Figure 1: A Five-Layer Functional Framework for the Epitome Intelligent Experimental Platform. (A) Functional Framework of Epitome. (B) Epitome supports a broad range of domain studies
The first layer, the Foundation Model Layer, anchors the platform with state-of-the-art LLMs and multimodal models, such as GPT-4, Gemini, and Deepseek-R1. It allows researchers to fine-tune generation parameters through the "My Bot" module, enhancing predictability and consistency in experimental outcomes. The platform also leverages Retrieval-Augmented Generation (RAG) technology to enable real-time access to external information, overcoming static response limitations and improving participant-AI interactions.The second layer accelerates the translation of experimental insights into deployable tools. Epitome offers a Workflow Auto-Planning Algorithm and integrates with Dify, a low-code development platform, enabling low-code and no-code development through an intuitive visual interface. Researchers can easily define prompts, contexts, and plug-ins, insert predefined functions, call specific LLMs or external APIs, and extract linguistic features and behavioral logs in real time. This layer lowers the technical threshold for social science researchers and enhances methodological transparency.The third layer forms the Human-AI Collaborative Experimental Environment, featuring three modules: "My Bot," "My Chatroom," and "Town Simulation." These modules support intelligent dialogue, multi-agent human-AI interactions, and fully configurable virtual environments populated by AI agents. Researchers can customize parameters, control experimental variables, and record detailed event logs, enabling unprecedented levels of experimental scale and ecological validity.The fourth layer operationalizes experimental rigor through multi-level randomization and versatile intervention channels. The "My Materials" module allows researchers to upload intervention materials, while randomized intervention assignment ensures a more robust experimental design. Additionally, Epitome’s open feature promotes knowledge sharing and enhances research transparency and reproducibility.The top layer provides a unified data-acquisition cockpit, integrating data visualization and collection. The "My Questionnaire" module supports diverse data-collection formats, and a real-time dashboard consolidates analytics panels for system configuration, topic and sentiment analysis, interactive text analysis, and basic statistics. This suite enables continuous monitoring, anomaly diagnosis, and on-the-fly parameter adjustments.Throughout the five layers, Epitome features a canvas-based interactive experimental design and an experiment management system, allowing researchers to flexibly assemble and execute complex procedures. The platform successfully replicates classic interdisciplinary studies, demonstrating its practical value in advancing the reproducibility of behavioral and social science experiments.
Traditional socioal science experiments are designed for human participants, which leads to many difficulties when incorporating LLMs(see in Table 1). Some specialized experimental platforms, such as PsyToolkit, involve precise time measurement and support the collection of human eye movement and EEG data. However, few platforms currently support the design of psychological experiments with complex human–computer interaction processes. Otree is a classic open-source experimental platform that supports complex multi-person interaction experiments. There have already been many open-source works based on Otree that attempt to integrate LLMs into the experimental process, but they often adopt a patch-like approach. For example, in (Engel et al., 2024), the depth of LLM integration into experiments is limited, and researchers also find it difficult to customize the agents. This makes it hard for LLMs to autonomously participate in complex human–computer interaction experiments as agents. Meanwhile, many practical agent-building platforms, such as Coze and Dify, have emerged in commercial scenarios. They support users in constructing complex agent workflows and offer rich external auxiliary modules. However, it is difficult for researchers to directly use them for rigorous psychological experimental research.The emerging Epitome platform deeply integrates psychological experimental paradigms with large language model (LLM) technology, opening up new methodological pathways for interdisciplinary research. Compared with existing psychological experimental platforms, Epitome has advantages in LLM integration. Unlike traditional survey tools such as Qualtrics, which lack a real-time interactive environment, and open-source psychological experimental platforms such as PsyToolkit and Cognition.run, which rely on programming skills, Epitome systematically solves the technical bottleneck of LLM integration into the experimental process through its canvas-based visual design system. Compared with some visualization experimental platforms that partially support LLM tools, such as Labvanced and Gorilla, Epitome supports basic model access and achieves deep integration of LLMs in the experimental environment through agent workflows and multimodal interaction frameworks. Compared with the limitations of group experimental platforms such as Otree, which rely on third-party toolkits to achieve LLM interaction (Engel et al., 2023; McKenna, 2023/2023), the natively integrated LLM system and multi-agent simulation system of Epitome can ensure the integrity and traceability of experimental data.
Figure 2: Epitome’s Functions Compared to Existing Social science Experiment Tools
By constructing $55$ layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com222www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
Its innovations are reflected in three principal dimensions: First, as an AI platform, Epitome integrates the advanced capabilities of Large Language Models (LLMs), fully leveraging their potential in the domains of AI for Social Science and Social Science of AI. Epitome’s modular, extensible architecture and flexible interfaces support diverse LLMs, simulations, and scenarios. This integration applies to experiment design, experimental environments, intervention strategies, and participant management. By harnessing its Foundation Model layer, Epitome offers diverse Human–AI Collaborative Experimental Environments, enabling researchers to adjust study parameters in real time. This flexibility broadens the methodological repertoire for AI-focused social science experiments. Second, as an experimental platform, Epitome embeds the principles of experimental science at the core of its digital design to promote "integrated experimentation." The platform’s architectural design is rooted in the "Intervention–Comparison–Outcome" principle, the fundamental logic of experimental science. To meet the core requirements of social science research, it features innovative designs such as group classification (supporting the creation of treatment and control groups) and an adaptive randomization mechanism that facilitates efficient hypothesis testing. Additionally, as a digital open platform, Epitome supports large-scale integrated experiments [35], mitigating limitations on experiment size and enhancing social validity. Third, as a research tool platform, Epitome adheres to the principle of "usability as a tool." Epitome simplifies research workflows with its Canvas-Based Interactive Design, merging the traditional four-phase, twelve-task process into a unified pipeline(see SI.1 Task) This end-to-end solution offers: a. Visual Experiment Management: Comprehensive visualization and real-time control of the entire research lifecycle.b. Intelligent Workflow Automation: Dynamic load balancing for scalable analysis and traceable data management.c.Low-Code Tools: Plug-and-play components that accelerate experiment design without coding. Accessible via dedicated portals, researchers can design studies through the Researcher Portal, while participants engage through a streamlined interface. The platform’s data-driven approach enables personalized experimental paradigms, systematic hypothesis testing, and enhanced comparability across studies. By integrating AI-driven complexity management, Epitome significantly improves research efficiency and methodological rigor in social science studies.
Figure 3: Visual representation of Case Study 1. (A)Case 1 examines the impact of generative artificial intelligence on managerial productivity. Forty-six managers were randomly assigned to either a treatment group, which utilized ChatGPT to complete two writing tasks, or a control group that completed the same tasks without AI assistance. The treatment group also provided feedback on time investment and self-efficacy. (B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 1). (C)To facilitate experimental replication in Case Study 1, Epitome embeds large language model dialogue capabilities within the survey, thereby enriching human–AI interaction and streamlining data collection.
How Epitome promote AI for SS and SS of AI
Epitome can support a wide range of AI for SS and SS of AI experimental researches across many disciplines.(see in Figure 1.B)
AI as Experimental Participants: Simulating Human-like Behavior and Controlled Experimental Design
The human-like intelligence of LLMs refers to their ability to mimic human behavior, which has led to their increasing application in social science research [36]. As the ability of large language models to simulate human-like characteristics continues to evolve, there has been a growing body of research utilizing LLMs as representations of human entities [37, 38, 39]. Epitome provides multidimensional support for such studies: Researchers can use Epitome’s My Bot module for personalized customization, effectively assisting in the creation of hypothetical scenarios [40]. Through prompt engineering and foundational model configuration, researchers can create My Bot simulations of groups with different demographic characteristics, such as personality traits, age, or cultural background differences, and showcase credible human behavior agents based on these varying traits.
The credible agents of human behavior realized by large models manifest as individual biases at the individual level and as consistency with human group behavior at the group level. From the perspective of individual behavior research, using economic experiments as an example, Epitome allows for the configuration of My Bot’s prompts to create "selfish" or "fair" chatbots. These can then be placed in a Chatroom scenario based on the Ultimatum Game, enabling the observation of decision-making biases of large models with different characteristics in the game [41]. From the perspective of group behavior consistency, Epitome supports the background settings of the Chatroom and sub-process configurations, reproducing classic experiments that span economics, psycholinguistics, and social psychology, such as the Ultimatum Game, Garden Path Sentences, and the Milgram Shock Experiment, allowing researchers to observe whether the results of simulated behavioral replication closely match real-world outcomes [42]. These customized large language models (LLMs) can support a variety of controlled experimental designs, including behavioral economics experiments [43], strategic game experiments [41], and social dilemma experiments [44].
Epitome also assists researchers in creating customizable "silicon samples" to support the simulation of human-like behavior. Large language models have demonstrated potential in providing preliminary psychological support and assisting with psychological therapy [45]. Epitome can be used to create virtual agents that simulate therapeutic conversations. By configuring prompts or workflows, researchers can provide different therapeutic methods and techniques as the basis for these conversations, helping them explore their effects and roles. At the same time, researchers can customize the personalities, habits, past experiences, and background scenarios of multiple roles through these modules, constructing fully large language model-driven simulated communities to observe human-like individual behaviors and emergent behaviors [10]. Virtual therapeutic agents created by My Bot interact with individuals experiencing mild stress, and through the Chatroom or Workflow modules, the interactions are set up and the process data of the interactions are recorded. Analysis of these interactions shows that the use of agents has a positive impact on reducing stress levels, but also highlights the importance of ethical issues such as privacy protection and informed consent. Additionally, the multi-agent virtual environment supported by Epitome, through the substitution of real human participants with multiple simulated agents, creates an effective social experiment sandbox that isolates the real world, thus avoiding the ethical issues that might arise in real human participant experiments, particularly reducing the moral risks associated with human subjects [46].
AI as Social Actors: Multi-Agent Interaction and Group Dynamics Modeling
The application of AI in simulating social behavior and group interactions is increasingly prevalent, offering new tools and perspectives for social science research. Epitome supports the creation of AI-driven social interaction scenarios, enabling researchers to construct complex multi-agent interaction environments to study group dynamics, cooperation, competition, and the emergence of social norms, as well as to explore the psychological responses of individuals and groups. Focus groups and joint decision-making are commonly used tools in social science research and are well-supported in Epitome. For example, by using the Chatroom module to simulate political debate scenarios, researchers can design sub-processes in the Chatroom according to political debate rules, set the background of the Chatroom based on the debate topic, facilitate the debate process, and record the process data. This allows researchers to investigate whether AI agents exhibit behaviors consistent with their self-interest and altruism in non-social decision-making tasks and dictator games [47].
Taking the signal game task of human-AI collaboration as an example: Researchers can predefine the type structure of the game (such as the role allocation between sender and receiver), cost parameters (such as the sending costs of different signals), and the payoff matrix as environmental variables. During the experiment, the agents autonomously read these environmental variables and, based on their built-in decision models, analyze the current game state in real-time. The system simultaneously computes the signal choices made by human participants and the strategy responses of the agents, and instantly updates the accumulated payoffs for both parties. The agent can also dynamically adjust its signal interpretation strategy based on historical interaction data, and even proactively send signals with strategic intent to the human participants. The interface will simultaneously display a summary of the agent’s "thought process," the historical selection records of both parties, and the real-time changes in payoffs. Additionally, researchers can set attitude scales at key decision points to evaluate the participants’ understanding and trust in the agent’s strategy, while behavior metrics such as decision-making time and strategy switching frequency can be recorded through data tracking. This allows for an in-depth analysis of the bidirectional mechanism of signal transmission and the joint strategy evolution process in true human-AI collaboration.
Human-AI Collaboration: Process Reengineering and Human-AI Synergy Effect Analysis
Epitome supports researchers in exploring collaboration models between AI and humans at the application level. For instance, the Workflow module effectively supports AI-assisted experimental design for generating scientific hypotheses [48]. Furthermore, by combining the My Model and My Bot modules, researchers can integrate third-party fine-tuned large language models through the My Model module and then configure the base model, prompts, and other parameters via My Bot, enabling the generation of psychological hypotheses based on fine-tuned large language models [49]. Research findings indicate that the hypotheses generated by the model are not merely replicas of previous human hypotheses; their clarity, impact, and originality are comparable to those generated by humans, with no significant differences.
Furthermore, researchers can optimize the Workflow module continuously in conjunction with various experimental intervention modules, expanding and iterating more complex experimental processes. For example, by adding a chatbot based on a fine-tuned large language model as the foundational model in the workflow [49], incorporating modules that set up stepwise questioning [50], and introducing adversarial dialogues [48], the quality of hypotheses can be enhanced. In combination with the My Questionnaire module, the application of the Workflow module can also call upon SOTA LLMs for data fitting [51], helping researchers improve analytical accuracy during the data collection process.
At the same time, Epitome supports the customization of interaction process and outcome data recording by setting environment variables, thereby facilitating the exploration of diverse research questions and experiments. In collaborative programming experiments, the Workflow module can record the dialogue rounds between human programmers and My Bot in real-time. For example, the platform can facilitate the analysis of cognitive load transfer patterns when AI assistance is involved, examining how the rate at which humans adopt AI-generated code suggestions changes with varying task complexity [52]. This collaborative model demonstrates the viability of the "tool-enhanced research" approach within the broader framework of AI applications in social science [36].
AI as a Social Observer and Analytical Tool: Large-Scale Data Analysis and Non-Reactive Research
Epitome supports large-scale data analysis by setting up and combining multiple AI agent workflows, effectively assisting researchers in qualitative data analysis, with results that can be validated by human researchers. Existing studies show that large language models can be effectively applied to four common text classification tasks in current social science research: sentiment analysis, stance detection, hate speech detection, and misinformation detection [36]. For instance, in sentiment analysis tasks, ChatGPT has performed exceptionally well in three text-based mental health classification tasks [53].
The My Bot and My Questionnaire modules in Epitome form the core toolchain for processing social data. Integrating a fine-tuned My Bot can perform sentiment analysis on interview texts. For example, in AI ethics research, researchers use AI to classify public comments based on ideology [54]. The My Questionnaire module supports AI-driven conversational surveys, providing conversational support and necessary assistance in progressing the survey by integrating a large language model into the questionnaire response page.
In non-reactive research, the My Bot module can support the processing of historical text corpora (e.g., newspaper articles from 1980-2020), using word vector analysis to identify shifts in social values. In more complex studies, researchers have designed a 50-billion-parameter language model for the financial sector—BloombergGPT [55]. Similar customized models can be integrated through the My Model module, then used as the foundational model for My Bot and role agents in the Chatroom, applied to specific experimental processes.
Human and Systematic evaluation of LLMs
Human evaluation is also a crucial method for exploring the interactive capabilities of LLMs . For example, when evaluating the cognitive abilities of large models, researchers can design prompts and adjust model parameters in My Bot to trigger specific cognitive biases and interact with participants, thereby revealing potential cognitive biases inherent in the LLM itself. Additionally, by embedding classic cognitive psychology scales through the My Questionnaire module, the platform supports comparative experiments between humans and AI across various cognitive tasks, such as perceptual judgment [56], logical reasoning [57], and decision-making behavior [58]. This enables the investigation of the cognitive rules demonstrated by AI and the differences between these rules and human cognitive architectures.
By constructing complex experimental workflows on the platform, the influence of LLM communication styles (such as framing effects and option order bias) on human decision-making can be systematically examined. In more complex experimental scenarios, such as multi-agent collaboration, researchers can achieve deep customization using the Dify Workflow development platform. For instance, in studies on investment decision biases, researchers can drag and drop to generate and connect logic modules for conditional judgments, large models, and iterations, set multi-agent interaction logic to present investment options through a gain/loss framework, and incorporate external data using the knowledge base module, completing application-level development.
Epitome’s Workflow development platform not only provides technical support for such experiments but also enables precise randomization design for experimental and control groups within a controlled environment, deeply embedding the "intervention-control-comparison" research paradigm into the data analysis process. Moreover, the platform’s multidimensional data tracking and collection mechanisms offer full-chain data support for researchers, facilitating the construction of a detailed human-AI decision-making bias map.
Case Study
In order to assess the robustness and versatility of the Epitome, we conduct a series of replications based on three influential studies [59, 60, 61]. We validated Epitome against established methodological standards in experimental social science through systematic comparison with traditional methods. Our validation approach assessed: (1) internal validity, by examining whether Epitome’s randomization procedures and blinding capabilities prevent confounding; (2) construct validity, by testing whether AI-human interactions in Epitome accurately represent theoretical constructs; and (3) external validity, by comparing results from Epitome experiments with findings from traditional laboratory studies. The three case studies demonstrate how Epitome satisfies these validation criteria while enabling new experimental possibilities beyond conventional methods.
Case study 1: Productivity Effects of Generative AI
To demonstrate Epitome’s versatility in facilitating complex human-AI social experiments, we replicated the recent study “Experimental Evidence on the Productivity Effects of Generative Artificial Intelligence” [60], published in Science. This study was replicated using the Epitome platform, conducting a repeated experiment with management professionals while maintaining alignment with the original study’s design and supporting materials (Figure 2.A). We recruited 46 participants who self-identified as experienced, college-educated managers (Dataset S1). Eleven participants were excluded from the analysis due to misreporting their occupation or failing to meet the specified time requirements in the exact-time group, resulting in a final sample of 35 participants. Participants were randomly assigned to either the treatment or control group. Each participant completed two incentivized writing tasks relevant to managerial roles, designed to simulate real-world responsibilities such as drafting performance evaluations and composing strategic communications. Between the two tasks, the treatment group was instructed to register for ChatGPT and received guidance on its use, whereas the control group was asked to register for the LaTeX editor Overleaf, ensuring comparable time and effort investment in learning a new tool. To ensure methodological rigor, all experimental procedures—including task instructions, incentive structures, evaluation criteria, and data collection protocols—were meticulously replicated from the original study(see Figs. S1–S7 for the detailed application of the experimental design).
The replication results were highly consistent with the original study. During the data analysis phase, Epitome’s data visualization tools effectively supported researchers in tracking experimental progress and analyzing participants’ demographic information and questionnaire responses. Preliminary findings showed that ChatGPT significantly reduced task completion time and improved work quality, with the treatment group outperforming the control group (see Fig. S8). Further analysis indicated that ChatGPT had a positive impact on participants’ work experience and self-efficacy. Participants reported significant improvements in task enjoyment, perceived skill, and efficiency, suggesting enhanced confidence and satisfaction (see Fig. S9). However, no significant changes were observed in participants’ concerns about AI’s impact on their careers, their excitement about AI, or their optimism regarding AI’s future in their field (see Fig. S10), indicating that short-term use of ChatGPT did not substantially alter their broader perceptions of AI and workplace automation.
Epitome provided a unique opportunity to capture more granular data on AI usage. By integrating real-time data tracking within the My Bot module, we were able to observe the number of interactions with the AI and the prompt content, revealing that fewer, more specific prompts led to better performance. This suggests diminishing returns when multiple AI-generated prompts were used, a nuance not observed in the original study. Additionally, Epitome’s automatic logging and randomized participant allocation ensured a higher level of experimental control, reducing human error in the setup and execution of the study. This case study demonstrates that Epitome can effectively replicate complex, multi-condition experiments in organizational settings, offering insights into AI’s role in enhancing productivity. Researchers in fields such as management, organizational behavior, and workplace productivity can leverage Epitome to explore the broader applications of generative AI in professional tasks. Furthermore, the platform’s ability to automate data collection and randomization makes it a powerful tool for future experimental research on AI-driven workplace interventions.
Figure 4: Visual representation of Case Study 2. (A)Case Study 2 explores the influence of generative AI on creative writing. Thirty-one writers composed stories with varying levels of AI support. Their work was evaluated through expert blind reviews as well as multidimensional assessments by lay evaluators, focusing on writing quality, originality, emotional expression, and perceived authorship. (B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 2). (C)The Dify-based low-code workflow supports the experimental design in Case Study 2, particularly by limiting the number of AI-generated ideas and streamlining the distribution of experimental materials.
Case study 2: Human Creativity and Generative AI
Figure 5: Visual representation of Case Study 3. (A)Case 3 investigates AI-driven role interaction in moral dilemma scenarios. Seven fictional characters were developed and introduced in the “AI Town” simulation to model AI-to-AI behavior. These character setting were then converted into prompts and deployed in a multi-agent chatroom, where additional AI agents participated in extended social interactions. The participants engaged in courtroom-style simulations, interacting with AI agents to explore moral decision-making and ethical dynamics in a controlled environment. (B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 3). (C)The AI Town multi-agent simulation environment and the multi-agent chatroom broaden research methodologies by enabling comparative studies of human versus AI decision-making in moral dilemmas for Case Study 3, thereby advancing social science theory.
In the second experiment, we replicate the study published in Science, titled “Generative AI enhances individual creativity but reduces the collective diversity of novel content” [61]. The purpose of the replication is to assess how Epitome can support high-fidelity replication of creative tasks and to extend the original findings by testing them in a multilingual setting (Chinese) and evaluating AI’s impact on both individual creativity and collective diversity of outputs.
The original experimental design involved three groups: a control group (Human-only), a group with one AI-generated suggestion, and a group with five AI-generated suggestions. We replicated the task structure, randomization process, and evaluation criteria. The only modification was the introduction of a Chinese language condition to assess the cross-cultural generalizability of the original study’s findings. The multilingual compatibility of Epitome effectively supports the broad recruitment of participants from diverse language backgrounds. Epitome’s ‘My Material’ and ‘My Bots’ modules were used to ensure that the AI interactions were faithfully replicated, and participants were randomly assigned to one of the three experimental conditions (see Figs. S11–S21 for the detailed application of the experimental design)..
In line with the original study, we found that generative AI significantly enhanced individual creativity, particularly for participants with lower baseline creativity(see Fig. S22 and Tables S8–9). Novelty increased by 10–12% and usefulness by 11%, with the greatest improvements observed in the low-creative group (see Fig. S23). However, we also observed a significant reduction in collective diversity, with AI-assisted outputs showing increased similarity (9–11%) compared to the control group(see Fig. S24). These results align with the original study’s findings and suggest that while AI can enhance individual creativity, it may also lead to a reduction in the diversity of collective creative outputs(see Supplementary Information SI.4 for additional experimental results).
Epitome provides valuable insights into the non-linear effects of AI input on creativity. Specifically, while a single AI suggestion improved output quality, five suggestions led to a reduction in diversity, confirming diminishing returns on AI assistance. The platform’s real-time tracking of AI interactions enabled detailed analysis of how AI-generated ideas influence creative work. This feature, along with the platform’s flexibility in adjusting the number and type of AI suggestions, offered a nuanced understanding of AI’s impact on creativity. In the replication study, participants were randomly assigned to one of three groups—no AI assistance, one AI suggestion, or five AI suggestions—and to one of three creative themes (jungle, planet, or ocean). This design, which required the deployment of multiple agents across different scenarios, significantly increased experimental complexity. To manage this, we utilized Epitome’s “My Workflow” module to create a multi-stage dialogue workflow, enabling precise processing of user inputs and generating targeted creative outputs. Epitome’s integrated capabilities not only streamlined the experimental process but also provided deeper insights into how varying levels of AI assistance shape creative outcomes, contributing significantly to our understanding of AI’s role in human creativity.
This case study demonstrates Epitome’s capability to replicate and extend creativity research in AI-driven environments. The platform’s flexibility in adapting experimental designs across multiple languages and cultural contexts makes it an ideal tool for cross-cultural studies on creativity. Future research could leverage Epitome to explore the role of AI in creative industries, such as advertising, design, and content creation, and examine how AI can be used to foster creativity while balancing the need for diversity in outputs.
Case study 3: Heinz Dilemma Experiment
To validate the role of Epitome in the design of social moral research experiments, we extended Kohlberg’s Heinz Dilemma experiment. The study was conducted in two phases: the first phase aimed to explore the decision-making processes of AI agents in moral dilemmas, while the second phase focused on how human participants interact with AI agents in a simulated environment. Through these two phases, we examined the thinking processes of both AI and human participants in ethical decision-making, particularly the conflict between justice and care ethics.
We utilized two Epitome modules for the experimental design. In the first phase, using Epitome’s AI Town module, we created an AI community based on the Heinz Dilemma framework and developed seven AI characters with distinct backgrounds, motivations, and objectives, based on roles from the original experiment. These AI characters were then placed within the community to make decisions reflecting moral judgments (see SI.6 for detailed role settings). In the second phase, we expanded the experiment into a specific scenario using Epitome’s multi-agent chatroom module, where LLM-based AI agents engaged in interactions with human participants(see SI.7 for detailed role settings). This phase incorporated a simulated trial scenario, inspired by the common law judicial process, in which participants were assigned roles as judges, jurors, and witnesses, each representing different perspectives on the ethical dilemma. Participants engaged in the trial process in the chatroom, debating with AI agents and performing a series of tasks, ultimately rendering a verdict on whether Heinz was guilty or not. Through Epitome’s real-time interactions and comprehensive logging, we collected valuable data that aids in analyzing the different strategies and preferences employed by humans and AI agents in handling complex moral issues(see Figs. S30–S35 for the detailed application of the experimental design).
The AI Town simulation showed that AI agents consistently favored justice, even when Heinz attempted to defy the law in the fourth simulation, with the narrative remaining focused on justice-related themes, as confirmed by the LDA analysis (see Table S15). In the courtroom simulation, AI judges closely aligned with human judges. However, differences emerged in the roles of jurors and witnesses. Female human participants were more inclined towards care ethics, while male participants and AI agents prioritized justice. AI’s emotional responses remained stable and positive in both simulations, with sentiment scores averaging around 7/10, indicating minimal fluctuation and sustained positivity (Figs. 36a-b). Throughout the four iterations, AI consistently favored justice. In the first three iterations, Heinz, persuaded by his wife, chose actions aligned with justice (see Table S14). Even in the fourth iteration, when Heinz resisted, the LDA analysis confirmed AI’s continued preference for justice. In the courtroom, both AI judges and human judges favored a “guilty but lenient” verdict 87.5% and 75% of the time, respectively (Figs. 36c). Female judges showed a stronger preference for leniency, with 100% of female participants choosing this option compared to 50% of male participants. When acting as jurors, AI and human participants both leaned toward justice, though female human participants showed more variability, balancing justice with care (Figs. 36d). These results highlight the contrast in moral reasoning: human participants, especially women, are more influenced by care ethics, while AI remains focused on justice and legal norms. This aligns with previous studies on gendered moral priorities and extends the framework of AI-assisted adjudication proposed by de Sousa et al. . Moreover, the emotional stability of AI’s decisions contrasts with human emotional sensitivity, underscoring the difference between human and AI decision-making.
Epitome’s capability to simulate real-time multi-agent interactions offers valuable insights into the dynamics of moral reasoning. By capturing detailed logs of AI interactions, we identified a consistent bias towards justice in the AI’s decision-making, contrasting with the more diverse responses exhibited by human participants. Furthermore, the platform’s ability to dynamically adjust the emotional states of AI agents introduces additional complexity to the simulation, enabling an investigation into how emotional factors influence moral judgments. This case study highlights Epitome’s potential for exploring ethical and moral decision-making within both AI and human contexts. The platform’s flexibility in simulating complex multi-agent scenarios positions it as an ideal tool for studying the role of AI in ethical decision-making. Future research can utilize Epitome to examine AI’s role in legal decisions, conflict resolution, and policy design, as well as the influence of various social factors and psychological traits on moral reasoning.
Discussion
The integration of large language models into social science experimentation advances theoretical development in three dimensions. First, Epitome enables researchers to test theories of human-AI interaction with unprecedented precision, providing empirical evidence regarding how cognitive, emotional and social processes manifest in human-AI contexts. Second, the platform facilitates comparative analysis between human-human and human-AI interactions, revealing both similarities and distinctions that refine existing social theories. Third, Epitome’s capacity to simulate multi-agent environments creates new possibilities for examining emergent social phenomena that traditional experimental methods cannot easily capture. These theoretical contributions extend beyond technical capabilities, addressing fundamental questions about technology’s role in human social systems and providing an empirical foundation for developing more robust social theories for the AI era.
Epitome’s innovation lies in its seamless fusion of advanced AI capabilities with traditional social science experimental methodologies. By embedding LLMs into a user-friendly, web-based interface, the platform democratizes access to AI tools for social scientists, overcoming the technical barriers that have historically limited their adoption. This integration allows researchers to design and execute complex experiments with unprecedented ease and precision, supporting a wide range of interaction modes and experimental paradigms. The platform’s drag-and-drop interface and real-time monitoring capabilities streamline the experimental process, from initial design to data collection and analysis. This flexibility is crucial for addressing the multifaceted nature of human-AI interactions and their societal implications. Moreover, Epitome’s support for diverse experimental frameworks, including multi-agent interactions, virtual environments, and randomized controlled trials (RCTs), positions it at the forefront of integrative experimental design. The platform’s ability to record detailed logs of AI behaviors and dialogues provides researchers with rich datasets for in-depth analysis, further enhancing the platform’s utility in both basic and applied research.
The success of Epitome in replicating and extending seminal studies in the field of human-AI interactions highlights its potential as a proof of concept (POC) platform for industry applications. As AI technologies continue to permeate various sectors, understanding their impact on human behavior and social dynamics becomes increasingly critical. Epitome’s capacity to simulate real-world scenarios and capture nuanced interactions between humans and AI agents offers valuable insights for industries seeking to develop human-centric AI solutions. In sectors such as healthcare, education, and customer service, Epitome could serve as a testing ground for evaluating the effectiveness of AI-driven interventions. For example, healthcare providers could use the platform to assess how AI chatbots influence patient engagement and treatment adherence. Similarly, educational institutions could explore the impact of AI tutors on student learning outcomes and motivation. By providing a controlled environment for experimentation, Epitome enables industries to identify optimal AI deployment strategies and mitigate potential risks.
Furthermore, Epitome’s open-source nature and support for diverse LLMs make it an attractive platform for collaborative research and development. Its modular design allows for easy integration with proprietary models and workflows, facilitating the customization of experiments to meet specific industry needs. This adaptability positions Epitome as a scalable solution for organizations looking to innovate and optimize their AI applications. The continued development of Epitome could focus on expanding its capabilities to support even more complex and dynamic human-AI interactions. Incorporating advanced physiological signal acquisition, such as EEG and eye-tracking, would enhance the platform’s ability to capture the subtleties of human cognitive and emotional responses. Additionally, integrating virtual reality (VR) and augmented reality (AR) technologies could create immersive experimental environments that more closely mimic real-world scenarios.
Materials and Methods
Detailed descriptions of materials and methods are available in SI Appendix  [^3] https://github.com/epitome-AISS/epitome.
Technical Implementations
In terms of technical capabilities, Epitome delivers four key advantages: first, its modular architecture and flexible interface definitions allow seamless integration of diverse large language models, extensible simulation environments, and custom experimental scenarios; second, by embedding an intelligent agent workflow engine with dynamic load-balancing, the platform ensures efficient resource allocation and robust performance under high-concurrency LLM inference; third, a personalized database and knowledge-base management system provides full traceability of experimental data and knowledge artifacts, underpinning rigorous social-science data analysis, reproducibility, and validation; and fourth, dual support for low-code configuration and operator “plug-and-play” development drastically lowers the barrier to experiment design and deployment, accelerating iteration cycles and fostering innovation in human–AI interaction studies.
The Epitome platform integrates the 12 tasks into a single platform based on the B/S architecture, providing a user-friendly web interface that allows both researchers and participants to easily log in and use the system. The front-end web page utilizes the React framework, an open-source JavaScript library focused on building reusable user interface components (Appendix X). React significantly enhances application performance through its virtual DOM technology, while its declarative programming paradigm enables developers to build complex user interfaces more efficiently and predictably. The component-based architecture of React further promotes code modularity and reuse, providing great flexibility in development.
For backend services, the platform adopts the Spring Boot framework, an open-source Java-based framework designed specifically for creating independent, production-level Spring applications. Spring Boot’s automatic configuration and modular design simplify the startup and operation processes of Spring applications. It also features an embedded web server and provides various production-ready features, including health checks and external configuration support, ensuring stability and maintainability. In the backend, the large model inference service uses the LLMs framework, which efficiently manages key-value memory in the attention mechanism via Paged Attention technology. This enables rapid execution of models using CUDA/HIP graph and supports tensor and pipeline parallelism for distributed inference, significantly boosting the inference speed of large models.
Four-Tier Framework for Epitome’s Ethical Safeguard Strategy
Epitome incorporates ethical safeguards aligned with principles of responsible AI research and human subjects protection. The platform implements a four-tier ethical framework: (1) informed consent mechanisms that clearly disclose when participants interact with AI agents; (2) data minimization protocols that collect only information essential to research questions; (3) transparency features that allow participants to review and understand AI-generated responses; and (4) oversight mechanisms that facilitate IRB review of experimental designs. These safeguards address the unique ethical challenges of human-AI experiments, particularly regarding participant autonomy and potential manipulation concerns.
Core functions and operation procedures
We provide a user manual for all experiment designers, offering detailed guidance on the functions and operational procedures of the Epitome platform (see SI.8-SI.11, Epitome User’s Guide). As the platform is currently in the beta testing phase, interested researchers who wish to use Epitome to support their experiments are encouraged to contact the authors for access.
Data, Materials, and Software Availability
Purpose: Provide the URL of the original dataset for easy access and use by other researchers, along with an open access statement.
ACKNOWLEDGMENTS
Please include your acknowledgments here, set in a single paragraph. Please do not include any acknowledgments in the Supporting Information, or anywhere else in the manuscript.

### v2

@@
 Epitome:Innovative Methodologies for Integrating AI and Social Science
-Layout and Key Functions By constructing $55$ layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com [^1] http://www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
+Layout and Key Functions By constructing 555 layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com [^1] http://www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
 Figure 1: A Five-Layer Functional Framework for the Epitome Intelligent Experimental Platform. (A) Functional Framework of Epitome. (B) Epitome supports a broad range of domain studies
@@
 Figure 2: Epitome’s Functions Compared to Existing Social science Experiment Tools
-By constructing $55$ layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com222www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
+By constructing 555 layers from Foundation Model, Complex Application Development, Human-AI iteration experimental environment, randomization and experimental intervention design and Data Visualization and Data Collection (Figure 1.A), Epitome (epitome-ai.com222www.epitome-ai.com) provides researchers with a comprehensive and efficient experimental platform(see Supplementary Information SI.2 for platform architecture).
 Its innovations are reflected in three principal dimensions: First, as an AI platform, Epitome integrates the advanced capabilities of Large Language Models (LLMs), fully leveraging their potential in the domains of AI for Social Science and Social Science of AI.
@@
 By integrating AI-driven complexity management, Epitome significantly improves research efficiency and methodological rigor in social science studies.
-Figure 3: Visual representation of Case Study 1.
-(A)Case 1 examines the impact of generative artificial intelligence on managerial productivity.
-Forty-six managers were randomly assigned to either a treatment group, which utilized ChatGPT to complete two writing tasks, or a control group that completed the same tasks without AI assistance.
-The treatment group also provided feedback on time investment and self-efficacy.
-(B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 1).
-(C)To facilitate experimental replication in Case Study 1, Epitome embeds large language model dialogue capabilities within the survey, thereby enriching human–AI interaction and streamlining data collection.
 How Epitome promote AI for SS and SS of AI
@@
 Case study 1: Productivity Effects of Generative AI
-To demonstrate Epitome’s versatility in facilitating complex human-AI social experiments, we replicated the recent study “Experimental Evidence on the Productivity Effects of Generative Artificial Intelligence” [60], published in Science.
-This study was replicated using the Epitome platform, conducting a repeated experiment with management professionals while maintaining alignment with the original study’s design and supporting materials (Figure 2.A).
+Figure 3: Visual representation of Case Study 1.
+(A)Case 1 examines the impact of generative artificial intelligence on managerial productivity.
+Forty-six managers were randomly assigned to either a treatment group, which utilized ChatGPT to complete two writing tasks, or a control group that completed the same tasks without AI assistance.
+The treatment group also provided feedback on time investment and self-efficacy.
+(B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 1).
+(C)To facilitate experimental replication in Case Study 1, Epitome embeds large language model dialogue capabilities within the survey, thereby enriching human–AI interaction and streamlining data collection.
+To demonstrate Epitome’s versatility in facilitating complex human-AI social experiments, we replicated the recent study "Experimental Evidence on the Productivity Effects of Generative Artificial Intelligence" [60], published in Science.
+This study was replicated using the Epitome platform, conducting a repeated experiment with management professionals while maintaining alignment with the original study’s design and supporting materials (Figure 3).
 We recruited 46 participants who self-identified as experienced, college-educated managers (Dataset S1).
@@
 Furthermore, the platform’s ability to automate data collection and randomization makes it a powerful tool for future experimental research on AI-driven workplace interventions.
+Case study 2: Human Creativity and Generative AI
 Figure 4: Visual representation of Case Study 2.
@@
 (C)The Dify-based low-code workflow supports the experimental design in Case Study 2, particularly by limiting the number of AI-generated ideas and streamlining the distribution of experimental materials.
-Case study 2: Human Creativity and Generative AI
-Figure 5: Visual representation of Case Study 3.
-(A)Case 3 investigates AI-driven role interaction in moral dilemma scenarios.
-Seven fictional characters were developed and introduced in the “AI Town” simulation to model AI-to-AI behavior.
-These character setting were then converted into prompts and deployed in a multi-agent chatroom, where additional AI agents participated in extended social interactions.
-The participants engaged in courtroom-style simulations, interacting with AI agents to explore moral decision-making and ethical dynamics in a controlled environment.
-(B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 3).
-(C)The AI Town multi-agent simulation environment and the multi-agent chatroom broaden research methodologies by enabling comparative studies of human versus AI decision-making in moral dilemmas for Case Study 3, thereby advancing social science theory.
-In the second experiment, we replicate the study published in Science, titled “Generative AI enhances individual creativity but reduces the collective diversity of novel content” [61].
+In the second experiment, we replicate the study published in Science, titled "Generative AI enhances individual creativity but reduces the collective diversity of novel content" [61].
 The purpose of the replication is to assess how Epitome can support high-fidelity replication of creative tasks and to extend the original findings by testing them in a multilingual setting (Chinese) and evaluating AI’s impact on both individual creativity and collective diversity of outputs.
@@
 The only modification was the introduction of a Chinese language condition to assess the cross-cultural generalizability of the original study’s findings.
-The multilingual compatibility of Epitome effectively supports the broad recruitment of participants from diverse language backgrounds.
+The multilingual compatibility of Epitome effectively supports the broad recruitment of participants from diverse language backgrounds(Figure 4).
 Epitome’s ‘My Material’ and ‘My Bots’ modules were used to ensure that the AI interactions were faithfully replicated, and participants were randomly assigned to one of the three experimental conditions (see Figs.
@@
 This design, which required the deployment of multiple agents across different scenarios, significantly increased experimental complexity.
-To manage this, we utilized Epitome’s “My Workflow” module to create a multi-stage dialogue workflow, enabling precise processing of user inputs and generating targeted creative outputs.
+To manage this, we utilized Epitome’s "My Workflow" module to create a multi-stage dialogue workflow, enabling precise processing of user inputs and generating targeted creative outputs.
 Epitome’s integrated capabilities not only streamlined the experimental process but also provided deeper insights into how varying levels of AI assistance shape creative outcomes, contributing significantly to our understanding of AI’s role in human creativity.
@@
 Case study 3: Heinz Dilemma Experiment
+Figure 5: Visual representation of Case Study 3.
+(A)Case 3 investigates AI-driven role interaction in moral dilemma scenarios.
+Seven fictional characters were developed and introduced in the "AI Town" simulation to model AI-to-AI behavior.
+These character setting were then converted into prompts and deployed in a multi-agent chatroom, where additional AI agents participated in extended social interactions.
+The participants engaged in courtroom-style simulations, interacting with AI agents to explore moral decision-making and ethical dynamics in a controlled environment.
+(B)Epitome-Based Interactive Experimental Design Canvas Interface (Case Study 3).
+(C)The AI Town multi-agent simulation environment and the multi-agent chatroom broaden research methodologies by enabling comparative studies of human versus AI decision-making in moral dilemmas for Case Study 3, thereby advancing social science theory.
 To validate the role of Epitome in the design of social moral research experiments, we extended Kohlberg’s Heinz Dilemma experiment.
@@
 Through these two phases, we examined the thinking processes of both AI and human participants in ethical decision-making, particularly the conflict between justice and care ethics.
-We utilized two Epitome modules for the experimental design.
+We utilized two Epitome modules for the experimental design(Figure 5).
 In the first phase, using Epitome’s AI Town module, we created an AI community based on the Heinz Dilemma framework and developed seven AI characters with distinct backgrounds, motivations, and objectives, based on roles from the original experiment.
@@
 Even in the fourth iteration, when Heinz resisted, the LDA analysis confirmed AI’s continued preference for justice.
-In the courtroom, both AI judges and human judges favored a “guilty but lenient” verdict 87.5% and 75% of the time, respectively (Figs.
+In the courtroom, both AI judges and human judges favored a "guilty but lenient" verdict 87.5% and 75% of the time, respectively (Figs.
 36c).
@@
 These results highlight the contrast in moral reasoning: human participants, especially women, are more influenced by care ethics, while AI remains focused on justice and legal norms.
-This aligns with previous studies on gendered moral priorities and extends the framework of AI-assisted adjudication proposed by de Sousa et al.
-.
+This aligns with previous studies on gendered moral priorities and extends the framework of AI-assisted adjudication proposed by de Sousa et al..
 Moreover, the emotional stability of AI’s decisions contrasts with human emotional sensitivity, underscoring the difference between human and AI decision-making.
@@
 Technical Implementations
-In terms of technical capabilities, Epitome delivers four key advantages: first, its modular architecture and flexible interface definitions allow seamless integration of diverse large language models, extensible simulation environments, and custom experimental scenarios; second, by embedding an intelligent agent workflow engine with dynamic load-balancing, the platform ensures efficient resource allocation and robust performance under high-concurrency LLM inference; third, a personalized database and knowledge-base management system provides full traceability of experimental data and knowledge artifacts, underpinning rigorous social-science data analysis, reproducibility, and validation; and fourth, dual support for low-code configuration and operator “plug-and-play” development drastically lowers the barrier to experiment design and deployment, accelerating iteration cycles and fostering innovation in human–AI interaction studies.
+In terms of technical capabilities, Epitome delivers four key advantages: first, its modular architecture and flexible interface definitions allow seamless integration of diverse large language models, extensible simulation environments, and custom experimental scenarios; second, by embedding an intelligent agent workflow engine with dynamic load-balancing, the platform ensures efficient resource allocation and robust performance under high-concurrency LLM inference; third, a personalized database and knowledge-base management system provides full traceability of experimental data and knowledge artifacts, underpinning rigorous social-science data analysis, reproducibility, and validation; and fourth, dual support for low-code configuration and operator "plug-and-play" development drastically lowers the barrier to experiment design and deployment, accelerating iteration cycles and fostering innovation in human–AI interaction studies.
 The Epitome platform integrates the 12 tasks into a single platform based on the B/S architecture, providing a user-friendly web interface that allows both researchers and participants to easily log in and use the system.
@@
 As the platform is currently in the beta testing phase, interested researchers who wish to use Epitome to support their experiments are encouraged to contact the authors for access.
-Data, Materials, and Software Availability
-Purpose: Provide the URL of the original dataset for easy access and use by other researchers, along with an open access statement.
-ACKNOWLEDGMENTS
-Please include your acknowledgments here, set in a single paragraph. Please do not include any acknowledgments in the Supporting Information, or anywhere else in the manuscript.